Complete each ordered pair so that it satisfies the given equation.
Y=3X+9: (0, ), ( ,24), (2, )

Answer by mangopeeler07(462)

y=3x+9 for (0, )

plug in 0 for x.

y=3(0)+9
y=0+9
y=9

(0,9)
-----------------------------------------------------------------------
y=3x+9 for ( ,24)

plug in 24 for y

24=3x+9
15=3x
5=x

(5,24)
-----------------------------------------------------------------------
y=3x+9 for (2, )

Plug in 2 for x

y=3(2)+9
y=6+9
y=15

(2,15)